If 10 be the Arithmetic mean of roots of a wuadratic equation and 5 be the geometric mean , then equation is 

Grade:12

4 Answers

Deepak Kumar Shringi
5 years ago
AM = a+b/2=10
a+b=20
ab=5
x2-20x+5=0
Tony
5 years ago
Quadratic equatio
Tony
5 years ago
Am is a+b /2  So a + b = 20
GM is square root of ab. So a×b = 25
So eq is x2-20x+25=0 because quad eq is written as x2-Sx+P=0
Manda Snehashish Reddy
5 years ago
AM is 10 which means a + b = 2(10)
GM is 5 which means ab = (5)2
Use the formula : x2 -(a+b) +ab = 0
                          x2-20x+25=0
DKS sir made a silly mistake